OpenStudy (anonymous):

When solving an equation like 2x + 10 = 2(x + 5), and, when simplified, you get x = x, 10 = 10, or 0 = 0, what is the solution for x?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

it means there is infinite solutions....no matter what number you sub in for x, it will still equal. example : 2x + 10 = 2(x + 5)......lets say x = 1 2(1) + 10 = 2(1 + 5) 2 + 10 = 2(6) 12 = 12 (correct) 2x + 10 = 2(x + 5).....lets say x = 2 2(2) + 10 = 2(2 + 5) 4 + 10 = 2(7) 14 = 14 (correct) no matter what number you sub in for x, it will always equal

OpenStudy (anonymous):

ok thanxs

OpenStudy (phi):

the solution is all real numbers if you had to write it down, you might write it this way \[ - \infty < x < \infty \]

OpenStudy (anonymous):

what does moderator mean

OpenStudy (anonymous):

@phi